OpenStudy (anonymous):

Find the equation for a parabola with its focus at (5, 0) and a directrix of x = -5.

OpenStudy (anonymous):

Read the link and you'll know how to do these from now on :-) http://www.mathwarehouse.com/quadratic/parabola/focus-and-directrix-of-parabola.php

OpenStudy (anonymous):

Better to teach a man to fish than to give him a fish, eh? It's also got a graphing tool for you it appears. How nice of them

OpenStudy (anonymous):

Also helpful: Vertex form of the equation of a parabola is given by y=a(x-h)\(^2\)+k where (h, k) are the coordinates of the vertex

OpenStudy (anonymous):

a = 1/4d or d = 1/4a

OpenStudy (anonymous):

If you're still stuck on something, ask away :-)